# Multi-Microphone Speech Emotion Recognition using the Hierarchical Token-semantic Audio Transformer Architecture

### v1

@@ Abstract
Most emotion recognition systems fail in real-life situations (“in the wild” scenarios) where the audio is contaminated by reverberation. Our study explores new methods to alleviate the performance degradation of Speech Emotion Recognition (SER) algorithms and develop a more robust system for adverse conditions. We propose processing multi-microphone signals to address these challenges and improve emotion classification accuracy. We adopt a state-of-the-art transformer model, the Hierarchical Token-semantic Audio Transformer (HTS-AT), to handle multi-channel audio inputs. We evaluate two strategies: averaging mel-spectrograms across channels and summing patch-embedded representations. Our multi-microphone model achieves superior performance compared to single-channel baselines when tested on real-world reverberant environments.
keywords:
speech emotion recognition, human-robot interaction
@@ 1 Introduction
Speech Emotion Recognition (SER) is widely studied in the literature. Most of the reported studies deal with clean speech data and do not consider additive noise and reverberant environments typical to real-life applications. Only a few studies address the influence of reverberation and noise on SER. In these studies, the reverberant data is artificially generated by convolving the clean utterances with Room Impulse Responsess, either simulated [1, 2] or recorded in a real environment with various reverberation levels [3, 4]. Noise may also be added to the reverberant signals.
A survey of Human-Robot Interaction (HRI) in a noisy environment can be found in [5, 6]. However, this survey does not address the SER task. Previous works have shown the significant challenges of detecting emotional speech in large and reverberant rooms. Reverberation can influence the speech signal and negatively affect the predicted results [7]. The literature on using multiple microphones for SER is very scarce. In [8], the robustness of Automatic Speech Recognition (ASR) systems to emotional speech in noisy conditions is addressed. A Binaural Emotional Speech Recognition (BESR) system is proposed, enabling the simultaneous acquisition of the speaker’s emotional state and transcribing the uttered speech signal.
Devices equipped with multi-microphones are widely available nowadays. Adding the spatial information may improve the performance of audio processing tasks, including SER. However, learning-based algorithms are challenged by differences between training and test conditions, specifically a change in the microphone array constellations, e.g., when the number of microphones in train and test conditions is different.
@@
In the current work, we extend the HTS-AT architecture to accommodate multi-channel inputs, thus enhancing robustness against reverberations. Since only a limited amount of data is available for our task, we resort to fine-tuning already-trained models. Moreover, our scheme can be fine-tuned with a certain number of microphones and tested with another microphone constellation, including microphone positions and the number of microphones. We evaluated the SER performance using three datasets, Ryerson audio-visual database of emotional speech and song (RAVDESS) [22], Interactive Emotional Dyadic Motion Capture (IEMOCAP) [23], and Crowd-sourced Emotional Multimodal Actors Dataset (CREMA-D) [24]. We used real-life RIRs from the Acoustic Characterisation of Environments (ACE) Challenge [25] to add reverberation to the speech utterances.
@@ 2 Problem Formulation
Let $x(t)$ denote the anechoic signal in the discrete-time domain. An $M$-microphone array captures this signal after propagating in the acoustic enclosure. The received microphone signals are then given by $y_{i}(t)=\{x*h_{i}\}(t),i=1,2,\ldots,M$, where $h_{i}(t)$ is the RIR from the source position to the position of microphone $i$. The aim of this work is to classify the emotion given the observations $y_{i}(t);i=1,\ldots,M$.
@@ 3 Proposed Model
The proposed SER model is based on the Swin-Transformer [26],111github.com/microsoft/Swin-Transformer, a state-of-the-art hierarchical ViT [21] architecture, used for a variety of computer vision tasks, that utilizes shifted windows to capture long-range image dependencies. We adopted and modified a Swin Transformer architecture, namely the Hierarchical Token-semantic Audio Transformer (HTS-AT) [27], which aims to improve the performance and scalability of audio tasks, such as the AudioSet dataset [28]. The HTS-AT model is designed to achieve the best performance by reducing the number of parameters, requiring fewer GPU resources, and less training time than the AST architecture [27]. In the current work, we propose further modifying the HTS-AT architecture to better suit the audio processing requirements. The main change is to adapt the model for multi-channel audio processing, expanding its usefulness beyond the original single-channel design. Two alternative multi-channel pre-processing strategies are examined. The first strategy applies a summation of patch tokens derived from mel-spectrograms. The second strategy applies averaging of mel-spectrograms.
@@ 3.1 Preprocessing and Input Features
We assume that the sampling rate of the audio signals is 16 kHz. Each microphone signal is first analyzed by a short-time Fourier transform (STFT), with a window size of 1024 samples and a hop size of 160 samples. The STFT bins are then aggregated to construct the mel-spectrograms. After fusing the multi-channel information into a single stream, the regular HTS-AT architecture is applied.
@@ 3.2 Architecture
We will now elaborate on the HTS-AT architecture. The standard transformer architecture requires extensive computational resources due to the unmodified input token sequence length across all layers. This includes maintaining a large global self-attention matrix and calculating outputs and gradients at each step. The HTS-AT architecture is introduced to address these challenges. The HTS-AT is designed for supporting multiple audio tasks, e.g., classification, sound event detection, and source localization. It introduces two key architectural optimizations: a hierarchical structure and a windowed attention mechanism. The input audio mel-spectrogram is split into localized patch tokens using a convolutional Patch-Embed CNN layer of kernel dimensions $P\times P$, in which the patches are ordered by time segment and frequency bin. Then, tokens propagate through a series of Swin Transformer encoder groups. At the terminus of each group, a Patch Merging layer reshapes the token sequence into its original 2D mel-spectrogram. This layer merges neighboring patches and then embeds them back into a latent space of reduced length. Consequently, the memory requirements decrease exponentially with depth. Within each Swin Transformer block, attention is restricted to non-overlapping $M\times M$ squares, partitioning the token sequence. Calculating self-attention within each window subset substantially reduces computational complexity relative to full global attention while capturing localized relationships. As patch size increases downstream, windows encapsulate larger temporal and frequency contexts. Finally, the HTS-AT incorporates a token-semantic CNN layer after the last transformer block. This layer refines the output by grouping tokens, thus capturing information about their time frames and frequency bins. Consequently, this enhances classification by exploiting token relationships.
@@ 3.3 Multi-Channel Methods
Figure 1: Scheme of patch tokens mel-spectrogram segments.
Figure 2: Scheme of average mel-spectrogram.
Inspired by the study in [29], we embraced a Patch-Embed scheme where each channel undergoes a shared embedding layer followed by summation. This maintains flexibility in the microphone numbers used during fine-tuning and testing without altering the core HTS-AT architecture. Additionally, summing Patch-Embed outputs consolidates inter-channel information, enabling more robust representations suitable for distortion and reverberation conditions.
@@
Patch-Embed Scheme: On the left side of Fig. 1, each of the $M$ channels is analyzed to generate mel-spectrograms, which are then concatenated along the channel depth. SpecAugment [30] is applied collectively across all $M$ mel-spectrograms. After this, each mel-spectrogram is reshaped to the dimensions of a $256\times 256$ image and passes through the shared Patch-Embed layer. Next, a summation operation is performed across all $M$-encoded channels, making this input suitable for the pre-trained HTS-AT. Finally, a single representation is derived, which is then sent to the feedforward mechanism of HTS-AT for emotion classification.
Average Mel-Spectrograms: The spectrogram averaging technique reduces reverberation effects by consolidating information from all channels. Since the model input is a mel-spectrogram image, averaging enhances frequency bias in the resulting image. This process happens at the start of the signal processing phase, before the Patch-Embed encoder layer. On the left side of Fig. 2, each channel is individually processed into a mel-spectrogram and then combined with other processed mel-spectrogram channels. Subsequently, averaging is performed based on the specified number of input channels, resulting in a transformation from a multi-channel to a single-channel representation. After this transformation, the mel-spectrogram is augmented and structured into an image. Following this, the image is sent to the Patch-Embed encoder layer, functioning as a feature tensor with dimensions of $4096\times 96$, and then progresses to the feedforward mechanism within the HTS-AT framework for emotion classification. It is worth noting that since the pre-trained model was specifically trained to handle a single channel, this characteristic enables the smooth integration of the mentioned representation into the model architecture. In contrast to the Patch-Embed Scheme approach, this method emphasizes preprocessing before patch embedding. As a result, it allows fine-tuning of a model with a predetermined number of microphones, making it easier to test with different microphone configurations. This feature provides flexibility and autonomy in the use of any number of microphones.
@@
Table 1: The results of single-microphone fine-tuned HTS-AT model compared with the method in [12] on clean (non-reverberant) datasets. We present the weighted average accuracy (in percentage) of 20 models that achieved the highest accuracy on the validation set and tested on the test set.
@@
4 Experimental Study
Datasets: Our study comprised three speech emotion recognition datasets, namely RAVDESS[22], IEMOCAP[23] and CREMA-D[24] datasets. The RAVDESS dataset comprises 24 actors, evenly distributed between male and female speakers, each uttering 60 English sentences. Hence, there are 1440 utterances expressing eight different emotions: ‘sad’, ‘happy’, ‘angry’, ‘calm’, ‘fearful’, ‘surprised’, ‘neutral’, and ‘disgust’. All utterances are transcribed in advance. Consequently, the emotions are more artificially expressed as compared to spontaneous conversation. In this dataset, we decided to union the emotions ‘neutral’ and ‘calm’ as representations of ‘neutral’. Therefore, we fine-tuned our model on seven classes instead of eight. A major drawback of the dataset is the small number of utterances. The IEMOCAP dataset comprises approximately 12 hours of speech and consists of conversations between two people that are either improvised or played according to a pre-determined transcript that was chosen to evoke different emotions. We fine-tuned our model on four classes: ‘happy’, ‘sad’, ‘angry’ and ‘neutral’. CREMA-D is a dataset of 7442 original clips from 91 actors comprising 48 male and 43 female actors. Speech utterances were selected from a set of 12 sentences. The sentences were presented using one of six different emotions ‘anger’, ‘disgust’, ‘fear’, ‘happy’, ‘neutral’ and ‘sad’.
For the multi-channel simulated RIR datasets, we first split the original three datasets into 80% of the data as a training set, 10% as a validation set, and 10% as a test set. Then, we used the gpuRIR Python code package222github.com/DavidDiazGuerra/gpuRIR to add reverberation and simulate multi-channel microphone signals (four channels) with reverberation time in the range of $T_{60}=200-800$ ms with different room sizes and randomized microphone locations. By doing so, we managed to enlarge the datasets. For RAVDESS, 6863 reverberant speech samples were generated for training, and 852 samples were used for validation. We follow the same procedure for IEMOCAP with 7356 train samples and 2107 validation samples and for CREMA-D with 5945 train samples and 1487 validation samples. Our objective encompasses evaluating our model in real-world reverberant environments. To achieve this, we conducted tests utilizing the ACE RIR database [25], which comprises seven distinct rooms characterized by varying dimensions and exhibiting diverse ranges of $T_{60}$ captured by a mobile phone equipped with three microphones. The model was fine-tuned with the synthesized RIRs and evaluated with the test sets of the various datasets convolved with the ACE RIRs.
@@ 4 Experimental Study
RAVDESS IEMOCAP CREMA-D
@@
Lecture Room 1 ($T_{60}=638$ ms) 77.3% 80.6% 81.3% 61.3% 67.0% 67.4% 63.2% 66.4% 67.4%
Lecture Room 2 ($T_{60}=1220$ ms) 77.3% 78.6% 82.0% 63.4% 66.1% 68.7% 65.4% 67.3% 66.0%
Lobby ($T_{60}=646$ ms) 78.6% 82.0% 84.0% 61.8% 64.0% 65.1% 64.2% 66.2% 66.8%
Meeting Room 1 ($T_{60}=437$ ms) 73.3% 80.6% 82.6% 62.9% 66.7% 68.7% 64.4% 65.7% 65.7%
Meeting Room 2 ($T_{60}=371$ ms) 82.0% 83.3% 85.3% 59.5% 66.3% 64.2% 65.6% 66.6% 66.8%
Office 1 ($T_{60}=332$ ms) 76.0% 83.3% 80.6% 63.6% 68.3% 67.2% 64.1% 68.8% 67.3%
Office 2 ($T_{60}=390$ ms) 78.6% 81.3% 82.6% 59.5% 64.7% 65.4% 62.6% 64.0% 64.1%

### v2

@@ Abstract
The performance of most emotion recognition systems degrades in real-life situations (“in the wild” scenarios) where the audio is contaminated by reverberation. Our study explores new methods to alleviate the performance degradation of Speech Emotion Recognition (SER) algorithms and develop a more robust system for adverse conditions. We propose processing multi-microphone signals to address these challenges and improve emotion classification accuracy. We adopt a state-of-the-art transformer model, the Hierarchical Token-semantic Audio Transformer (HTS-AT), to handle multi-channel audio inputs. We evaluate two strategies: averaging mel-spectrograms across channels and summing patch-embedded representations. Our multi-microphone model achieves superior performance compared to single-channel baselines when tested on real-world reverberant environments.
Index Terms—  speech emotion recognition, human-robot interaction
@@ 1 Introduction
SER is widely studied in the literature. Most of the reported studies deal with clean speech data and do not consider additive noise and reverberant environments typical to real-life applications. Only a few studies address the influence of reverberation and noise on SER. In these studies, the reverberant data is artificially generated by convolving the clean utterances with Room Impulse Responsess, either simulated [1, 2] or recorded in a real environment with various reverberation levels [3, 4]. Noise may also be added to the reverberant signals.
A survey of Human-Robot Interaction (HRI) in a noisy environment can be found in [5, 6]. However, this survey does not address the SER task. Previous works have shown the significant challenges of detecting emotional speech in large and reverberant rooms. Reverberation can influence the speech signal and negatively affect the predicted results [7]. The literature on using multiple microphones for SER is very scarce. In [8], the robustness of Automatic Speech Recognition (ASR) systems to emotional speech in noisy conditions is addressed. A Binaural Emotional Speech Recognition (BESR) system is proposed, enabling the simultaneous acquisition of the speaker’s emotional state and transcribing the uttered speech signal. Devices equipped with multi-microphones are widely available nowadays. Adding the spatial information may improve the performance of audio processing tasks, including SER. However, learning-based algorithms are challenged by differences between training and test conditions, specifically a change in the microphone array constellations, e.g., when the number of microphones in train and test conditions is different.
@@
In the current work, we extend the HTS-AT [22] architecture to accommodate multi-channel inputs, thus enhancing robustness against reverberations. Since only a limited amount of data is available for our task, we resort to fine-tuning already-trained models. Moreover, our scheme can be fine-tuned with a certain number of microphones and tested with another microphone constellation, including microphone positions and the number of microphones. We evaluated the SER performance using three datasets, Ryerson audio-visual database of emotional speech and song (RAVDESS) [23], Interactive Emotional Dyadic Motion Capture (IEMOCAP) [24], and Crowd-sourced Emotional Multimodal Actors Dataset (CREMA-D) [25]. We used real-life RIRs from the Acoustic Characterisation of Environments (ACE) Challenge [26] to add reverberation to the speech utterances.
@@ 2 Problem Formulation
Let $x(t)$ denote the anechoic signal in the discrete-time domain. An $M$-microphone array captures this signal after propagating in the acoustic enclosure. The received microphone signals are then given by $y_{i}(t)=\{x*h_{i}\}(t),i=1,2,\ldots,M$, where $h_{i}(t)$ is the RIR from the source position to the position of microphone $i$. This work aims to classify the emotion given the observations $y_{i}(t);\;i=1,\ldots,M$.
@@ 3 Proposed Model
The proposed SER model is based on the Swin-Transformer [27],111github.com/microsoft/Swin-Transformer a state-of-the-art hierarchical ViT [21] architecture, used for a variety of computer vision tasks, that utilizes shifted windows to capture long-range image dependencies. We adopted and modified a Swin Transformer architecture, namely the Hierarchical Token-semantic Audio Transformer (HTS-AT) [22], which aims to improve the performance and scalability of audio tasks, such as the AudioSet dataset [28]. The HTS-AT model is designed to achieve the best performance by reducing the number of parameters, requiring fewer GPU resources, and less training time than the AST architecture [22]. In the current work, we propose modifying the HTS-AT architecture to suit the audio processing requirements better. The main change is to adapt the model to the multi-channel audio processing task, expanding its usefulness beyond the original single-channel design. Two alternative multi-channel pre-processing strategies are examined. The first strategy applies a summation of patch tokens derived from mel-spectrograms, and the second strategy applies averaging of mel-spectrograms. Our proposed approach first fuses the multi-channel information into a single stream. Then, the regular HTS-AT architecture is applied.
@@ 3.1 Preprocessing and Input Features
We assume that the sampling rate of the audio signals is 16 kHz. Each microphone signal is first analyzed by a short-time Fourier transform (STFT), with a window size of 1024 samples and a hop size of 160 samples. The log-absolute value of the STFT bins are then aggregated to construct the mel-spectrograms.
@@ 3.2 Architecture
We will now elaborate on the HTS-AT architecture. The standard transformer architecture requires extensive computational resources due to the unmodified input token sequence length across all layers. This includes maintaining a large global self-attention matrix and calculating outputs and gradients at each step. The HTS-AT architecture is introduced to address these challenges. The HTS-AT is designed for supporting multiple audio tasks, e.g., classification, sound event detection, and source localization. It introduces two critical architectural optimizations: a hierarchical structure and a windowed attention mechanism. The input audio mel-spectrogram is split into localized patch tokens using a convolutional Patch-Embed CNN layer of kernel dimensions $P\times P$, in which the patches are ordered by time segment and frequency bin. Then, tokens propagate through a series of Swin Transformer encoder groups. At the terminus of each group, a Patch Merging layer reshapes the token sequence into its original 2D mel-spectrogram. This layer merges neighboring patches and then embeds them back into a latent space of reduced length. Consequently, the memory requirements decrease exponentially with depth. Within each Swin Transformer block, attention is restricted to non-overlapping $M\times M$ squares, partitioning the token sequence. Calculating self-attention within each window subset substantially reduces computational complexity relative to full global attention while capturing localized relationships. As patch size increases downstream, windows encapsulate larger temporal and frequency contexts. Finally, the HTS-AT incorporates a token-semantic CNN layer after the last transformer block. This layer refines the output by grouping tokens, thus capturing information about their time frames and frequency bins. Consequently, this enhances classification by exploiting token relationships.
@@ 3.3 Multi-Channel Methods
Fig. 1: Scheme of patch tokens mel-spectrogram segments.
Fig. 2: Scheme of average mel-spectrogram.
@@
Patch-Embed Scheme: Inspired by the study in [29], we embraced a Patch-Embed scheme where each channel undergoes a shared embedding layer followed by summation. This maintains flexibility in the microphone numbers used during fine-tuning and testing without altering the core HTS-AT architecture. On the left side of Fig. 1, each of the $M$ channels is analyzed to generate mel-spectrograms, which are then concatenated along the channel depth. SpecAugment [30] is applied collectively across all $M$ mel-spectrograms. Subsequently, each mel-spectrogram is reshaped to the dimensions of a $256\times 256$ image and passes through the shared Patch-Embed layer. Next, a summation operation is performed across all $M$-encoded channels, making this input suitable for the pre-trained HTS-AT. The summation consolidates inter-channel information, enabling more robust representations suitable for reverberant conditions. The summed embeddings are then fed to the HTS-AT feedforward block for the actual emotion classification.
Average Mel-Spectrograms: On the left side of Fig. 2, the mel-spectrogram of each channel is calculated, followed by an averaging stage that produces a single-channel representation, regardless of the original number of channels. Since reverberation affects each channel differently, the averaged spectrogram will tend to be less reverberant. The averaged mel-spectrogram is then augmented and structured into an image. Subsequently, the image is fed to the Patch-Embed encoder layer, generating a feature tensor with dimensions of $4096\times 96$. Similarly to the first architecture, the HTS-AT feedforward block implements the emotion classification.
4 Experimental Study
In this section, we describe the experimental setup and present a comparative study of the proposed scheme and a baseline method.
4.1 Datasets
Our study comprised three speech emotion recognition datasets, namely RAVDESS[23], IEMOCAP[24] and CREMA-D[25] datasets.
The RAVDESS dataset comprises 24 actors, evenly distributed between male and female speakers, each uttering 60 English sentences. Hence, there are 1440 utterances expressing eight different emotions: ‘sad’, ‘happy’, ‘angry’, ‘calm’, ‘fearful’, ‘surprised’, ‘neutral’, and ‘disgust’. All utterances are transcribed in advance. Consequently, emotions are more artificially expressed than in spontaneous conversation. In this dataset, we decided to merge the emotions ‘neutral’ and ‘calm’ as representations of ‘neutral’. Therefore, the output of our model is on seven classes instead of eight. A significant drawback of the dataset is its small number of utterances.
The IEMOCAP dataset comprises approximately 12 hours of speech expressing four emotions: ‘happy’, ‘sad’, ‘angry’ and ‘neutral’. It consists of conversations between two people that are either improvised or played according to a pre-determined transcript chosen to evoke different emotions.
The CREMA-D is a dataset of 7442 original clips from 91 actors comprising 48 male and 43 female actors. Speech utterances were selected from a set of 12 sentences. The sentences were presented using one of six different emotions ‘anger’, ‘disgust’, ‘fear’, ‘happy’, ‘neutral’ and ‘sad’.
For the multi-channel experiments, we first split the original three datasets as follows: 80% as a training set, 10% as a validation set, and 10% as a test set. We note that the same actor may appear in several splits but not the same utterance.
We used the gpuRIR Python package222github.com/DavidDiazGuerra/gpuRIR to simulate reverberant multi-channel microphone signals (four channels) with reverberation time in the range of $T_{60}=200-800$ ms with different room sizes and randomized microphone locations. The room dimensions were uniformly distributed between 3 and 8 meters, with an aspect ratio ranging from 1 to 1.6. The room’s height was fixed at 2.9 meters. The sound source was placed at an elevation of 1.75 meters, while the microphones were positioned at 1.6 meters. Both the sound source and microphones were kept at least 0.5 meters away from the surrounding walls. By doing so, we were also able to enlarge the datasets. For RAVDESS, 6863 reverberant speech samples were generated for training, and 852 samples were used for validation. We follow the same procedure for IEMOCAP with 7356 train samples and 2107 validation samples and for CREMA-D with 5945 train samples and 1487 validation samples. Our objective encompasses evaluating our model in real-world reverberant environments. To achieve this, we conducted tests utilizing the ACE RIR database [26], which comprises seven distinct rooms characterized by varying dimensions and exhibiting diverse ranges of $T_{60}$. We used the subset captured by a mobile phone equipped with three microphones, which are ubiquitous and often used for speech interaction, making them a practical choice for real-world SER applications. The model was fine-tuned with the synthesized RIRs and evaluated with the test sets of the various datasets convolved with the ACE RIRs.
@@ 4.1 Datasets
Table 1: The results for the non-reverberant single-microphone case. The fine-tuned HTS-AT model is compared with [12]. We present the weighted average accuracy (in percent) of 20 models that achieved the highest accuracy on the validation set and tested on the test set.
@@
Room ($T_{60}$ [ms]) RAVDESS IEMOCAP CREMA-D
@@
Lecture 1 ($638$) 77.3 (70.6-84.0) 80.6 (74.6-86.6) 81.3 (74.6-87.3) 61.3 (57.1-65.2) 67.0 (63.0-71.0) 67.4 (63.6-71.2) 63.2 (60.1-66.6) 66.4 (63.2-70.1) 67.4 (64.2-71.0)
Lecture 2 ($1220$) 77.3 (70.6-83.3) 78.6 (71.3-84.6) 82.0 (76.0-88.0) 63.4 (59.5-67.4) 66.1 (62.4-70.0) 68.7 (65.0-72.4) 65.4 (62.0-68.9) 67.3 (64.0-70.8) 66.0 (62.5-69.3)
Lobby ($646$) 78.6 (72.0-84.6) 82.0 (76.0-88.0) 84.0 (78.0-89.3) 61.8 (58.0-65.8) 64.0 (60.0-67.8) 65.1 (61.3-68.8) 64.2 (60.9-67.8) 66.2 (62.8-69.8) 66.8 (63.3-70.2)
Meeting 1 ($437$) 73.3 (66.6-80.0) 80.6 (74.0-86.6) 82.6 (76.6-88.0) 62.9 (58.9-66.9) 66.7 (62.9-70.6) 68.7 (65.1-72.4) 64.4 (61.0-67.8) 65.7 (62.5-69.3) 65.7 (62.2-69.0)
Meeting 2 ($371$) 82.0 (76.0-87.3) 83.3 (76.6-89.3) 85.3 (80.0-90.6) 59.5 (55.7-63.6) 66.3 (62.4-70.3) 64.2 (60.2-68.1) 65.6 (62.4-69.0) 66.6 (63.4-69.8) 66.8 (63.6-70.2)
Office 1 ($332$) 76.0 (69.3-82.6) 83.3 (77.3-88.6) 80.6 (74.0-86.6) 63.6 (59.7-67.6) 68.3 (64.7-72.3) 67.2 (63.6-71.0) 64.1 (60.1-67.6) 68.8 (65.4-72.4) 67.3 (64.0-71.0)
Office 2 ($390$) 78.6 (72.6-84.6) 81.3 (75.3-87.3) 82.6 (76.0-88.0) 59.5 (55.5-63.8) 64.7 (61.0-68.8) 65.4 (61.5-69.6) 62.6 (59.4-66.1) 64.0 (60.6-67.6) 64.1 (60.6-67.7)
4.2 Algorithm Setup